January 5, 2012 The astigmatism problem is repaired. I’ve begun an overnight process that should improve image quality. The scope will be back online tomorrow morning.
January 5, 2012 There is a particle stuck on the objective lens that is creating uncorrectable astigmatism. The FESEM is unusable. I will attempt a repair today, so I have blocked the calendar for the entire day.
I’ll send you a status update later today.
November 27, 2011 The fine focus control is now repaired.
November 23, 2011 I’ve received a report that the EDS system on the FESEM is not working properly. I’ll fix it next week after the holiday weekend.
November 15, 2011 I am still working on the fine focus control. Hitachi sent the wrong part. I hope for a new one to show up this week. In the mean time, I found that the fine focus on the Windows slider bar works. Bottom left of screen.
November 9, 2011 The fine focus knob is not working, so Owen is in the process of fixing this part. In the mean time, you may still reserve and utilize the microscope, but you will not be able to adjust with the fine focus knob.
September 2, 2011 FESEM is up – During the repair I had to reset all the electronic alignments. I’ve realigned the scope for most popular kV settings, but you may find the need to make minor adjustments. Let me know if you have a problem. Don’t get in trouble by making changes you don’t have experience with.
September 2, 2011 FESEM down – I’m working on it now. Look for an update soon.
August 15, 2011 Have any of you found the Mac computer locked up recently? If so, when? I think I fixed it.
July 29, 2011 FESEM EDS computer problem: I’m having a problem with the EDS Mac computer freezing up intermittently. I can’t be sure if it it the computer or a program(s). I need your help to test this. Please close the Revolution EDS software when you finish your session.
Let me know if you find the system frozen. I’f you do find it frozen and can’t find me, depress the lighted button on the front of the CPU until the LED goes off and shuts the computer off. Wait 10 seconds and depress the button again to restart the system.
June 2, 2011 The computer logbook system is back online and I’ll be removing the paper logbook.
March 31, 2011 The computer logbook system on the FESEM is down. Please use the paper logbook next to the computer.
March 19, 2011 FESEM maintenance next week. The Hitachi engineers will be here next week to service the FESEM so there will be a few days when it is not available for use. Plan accordingly.
March 19, 2011 FESEM is working now. It was a heat problem.
March 19, 2011 The FESEM is unusable – no image. I have a call into Hitachi for help. Look for updates.
March 3, 2011 FESEM is working and can now be used.
March 3, 2011 The FESEM is down. I’m in touch with the Hitachi engineers.
January 5, 2011 EDS systems are working. Sorry for the delay.
December 20, 2010 The EDS system on FESEM, 6400 SEM and TEM will be offline beginning tomorrow afternoon. Offline means they won’t work and cannot be used.
I will bring them back online Monday, Jan 3rd.
December 13, 2010 We have installed a manipulator system inside the FESEM. A special sensor is attached to the device that can be struck and damaged when tilting the stage. The manipulator is modular and can be removed from the FESEM chamber and when it is removed, there is no chance to damage any part of it. My plan is to organize manipulator use for some periods and at any other time it will be removed from the FESEM. During those periods, I will restrict use of the scope to that kind of work. I will let you know in advance when we will schedule manipulator use.
November 30, 2010 The FESEM will be unavailable next week as we install an accessory. Please get your work done this week.
November 3, 2010 The accessory installation has been postponed until December. I’ve opened up the calendar for this Friday and then Monday through Wednesday next week.
October 29, 2010 FESEM EDS is working now
October 29, 2010 Note that the FESEM will be unavailable for 3 days next week and one day the week after. Get your work done now.
October 20, 2010 I found a broken power supply in the EDS computer. I’ve ordered a new one and it will be here Monday. I’ll install it right away so the system should be back online next Tuesday.
October 19, 2010 FESEM EDS is down. The computer is not working. I’ll check it tomorrow.
October 14, 2010 The FESEM will be offline during the entire week of Nov 1 for installation of an accessory. Please plan your work accordingly.
September 13, 2010 I am putting together a FESEM training session. To qualify, you must have completed the SEM course lab (MY4201 or MY5200) and have passed the practical exam.
The training sessions are 5-6 hours on consecutive days. I limit attendance to 4. The training will be help in early October.
September 13, 2010 The computer was misbehaving this morning so I’m running a cleanup routine that is taking longer than I’d expected. Call me before you come over to make sure the machine is back running.
July 20, 2010 I will begin offering the 6 hr training seminar for FESEM certification soon. Please let me know if you are interested in this training. You must have certification to use the JEOL 6400 SEM to take this training. I also ask that you take this training immediately before you plan to use the FESEM. Students that wait 6 months or more often forget what they learned and need to be retrained. This first training is free but we charge for a second training session. I will schedule training in small groups of 3-4.
Let me know if you want training.
June 21, 2010 the FESEM is operational and ready for use.
June 20, 2010 The FESEM is down. I’ll be back from vacation Tuesday and will check it then.
May 20, 2010 The FESEM is now available from 9am to 1pm tomorrow, Friday the 21st.
May 19, 2010 The Hitachi engineers will be working on the FESEM this afternoon. Call me at 369-1875 before you come for an appointment
May 3, 2010 You can schedule time on the FESEM now. The Hitachi engineer is here and want s to see if the scope HV/gun valve issue will develop again.
If we have trouble I will take it offline again.
May 1, 2010 There is an internal power supply problem in the FESEM and I’ve shut the system down. The Hitachi engineer will be here Monday to get us back online.
April 29, 2010 We have experienced an intermittent problem with the gun isolation valve. At times the valve will not respond until many seconds after turning off the HV. In some cases it will not open (or close) at all, and the HV cannot be turned on or off. When the system is not working the stage won’t go to the home position and the sample cannot be removed from the chamber.
I have tried a few fixes this morning and the valve seems to be working okay. If you see any problems with the valve, let me know right away. I need to test the system in a “fail” mode. My cell is 369-1875.
The Hitachi engineer is planning to come Monday.
April 28, 2010 Again today the gun valve did not open at the proper time. An error message appears on the computer screen. I got it to work but we need to watch it carefully. Let me know if you see anymore problems.
April 27, 2010 There is a report that the gun valve may not be closing when the HV is turned off. You should hear a loud hiss-thunk when the HV is turned off. If not, let me know right away.
April 7, 2010 I made a repair and tested it for 4 hours. It is okay to use it now. Please let me know if you see any error messages. Make a note of the message.
April 7, 2010 The FESEM is down (not working) due to a cooling water problem. I’m working on it.
April 6, 2010 The FESEM is out of order today.
March 28, 2010 I will vent the FESEM chamber today at 5 pm. The vacuum won’t be back to acceptable levels until tomorrow morning.
March 19, 2010 FESEM file saving problem is fixed now. You should be safe saving files.
March 16, 2010 I have not heard from Hitachi yet and I won’t be here tomorrow. In the meantime, you are free to use the scope. I suggest that you save a few images at a time, rather than wait until you’ve acquire 10 images. That may prevent loss of data. There is a problem saving image files on the FESEM. I’m waiting for a call from Hitachi about a fix. Don’t use it int he meantime.
January 21, 2010 We have installed 4pi hardware and Revolution software on the FESEM, 6400 SEM and TEM. Instructions for using the software can be found herehttp://electronmicroscopy.org/ Navigate to Hitachi S-4700 FE-SEM > Microanalysis. Note that the instructions are written for the FESEM and some microscope operations will be different. We are actively working on revised instructions for the 6400 and TEM. They should be published within a week or so. Few of you use standards-based quantitative microanalysis and instead use semi-quantitative analysis. I’ve mentioned in earlier messages that depending on the geometry and composition of your sample, error in the results could approach 20%. I think it would be a good idea to use an off-line microanalysis software package to verify your results. Please download that software, DTSA II, here. http://www.cstl.nist.gov/div837/837.02/epq/dtsa2/index.html This is Java software and like ImageJ is free (actually, you paid for it through federal taxes). Let me know if you have questions.
January 7, 2010 The 6400 SEM is working with new 4 pi software for EDS (same as FESEM and TEM) and a new revision of dPict (image acquisition). If possible, wait until next week to use it when I can show you how.
December 29, 2009 The FESEm is ready for use. The bake is completed and I aligned the column. The deflectors (Column Alignment window) are adjusted for 1kV. You will notice that those adjustments are farther out of alignment than usual. Let me know if you need help bringing them into alignment for your working conditions. I’d be interested to know if you find that the imaging performance is improved.
December 28, 2009 I am baking the FESEM gun overnight. It will be back online before noon tomorrow. The 6400 4pi EDS is calibrated and working. The 4pi x-ray mapping is working but needs calibration. The Geller imaging is working but needs calibration. I will complete both tasks this week. The WDS is not working and I will get to it as I have time. I am still conditioning the 4000 to 300kV. I’m making progress but there are 2 issues. First, I am having trouble getting the 300kV alignment straight. Second, the beam is flickering due to outgassing in the gun chamber. I noticed flickering at 200kV today too, but the scope was usable. I am returning the scope to 300kV and will run it at saturation until tomorrow morning when I have a 200kV job from 9 to noon. Again, if you have TEM work make an appointment with me so I can have the scope aligned at 200kV for you.
December 4, 2009 The battery backup is beeping in the FESEM lab. It is not a problem with the scope. I am repairing it now.
September 9, 2009 I’m looking for people who need (not want) FESEM training. You must have already completed the SEM course – you don’t qualify to take the training if you are in the SEM course now. Please do not sign up if you plan to use the FESEM at some future time. I will be offering these sessions as needed.
August 14, 2009 The FESEM EDS is working.
August 14, 2009 The FESEM EDS system is now up and running.
August 14, 2009 I need to adjust the EDS processor settings on the FESEM EDS system. That should go quickly. I’ll be working on it at 11 this morning.
June 18, 2009 The fine focus control quit working. The Hitachi engineer found a loose connector and the scope has been serviced and is ready to use.
June 8, 2009 I have completed the install and configuration of a new version of the Revolution EDS software for the FESEM. The user interface looks very much like the old software versions and the basic file management system is the same. Contextual menus are used to a great degree, accessed by a right mouse click. Besides the basic qualitative EDS function a few features are easier to use including x-ray mapping, linescans, digital imaging. Keep in mind that this is a new software version – not hardware – so we have not extended the system’s fundamental capabilities. We are building a eTraining module to assist you and in the meantime we’d be happy to take you through the software use.
June 1, 2009 I’ve installed new software on the EDS system. All is well, except the deat time is too high. I am pretty sure I know what it is and will make the appropriate adjustment.
February 11, 2009 We’ve identified the problem with the HV shutting off and I’ve ordered the parts – they should be here next week. In the meantime, I’ve bypassed the problem and the scope is usable. I doubt the the HV problem will resurface. Let me know if you observe anything out of the ordinary.
February 2, 2009 The HV continues to shut off unexpectedly. I am not sure why and neither is Hitachi. The Hitachi engineer is tentatively set to come next week.
Please continue to use the FESEM in the meantime and notify me immediately if the HV goes off. I am especially interested in a “rattling” noise that occurs with the HV shut off. If I am able to get to the scope before the noise stops, I may be able to follow the noise to the problem.
The FESEM maintenance scheduled for today, tomorrow and Wed is cancelled. I opened the calendar for your scope use.
January 30, 2009 I made a few repairs and the scope ran at 20kV all afternoon. I think it is ready for use. Let me know if you have any problems.
January 29, 2009 I may have identified and fixed the HV shut off problem today. I used the scope for 2 hours myself with no shutoff. Keep using it and let me know if it happens to you.
January 28, 2009 The HV is intermittently shutting off on the FESEM. I’ve got a call into to Hitachi. In the meantime, I’d say it’s okay to keep using the scope, but you may have trouble occasionally. I’d appreciate it if you listen and look for any odd things that happen when the HV shuts off. Especially keep an eye on the IP vacuum.
I’ll keep you updated.
The Hitachi engineer will come late next Monday and stay at least till Wednesday to look at some problems with the FESEM.
January 27, 2009 Please save images to the “D” drive only.
January 19, 2009 The FESEM is back online and ready to use. I ask that you let me know if you see anything unusual, especially with the vacuum system. Thanks for your patience.
January 12, 2009 The compressor is dead on the chiller. We will order a new one, installation is tentatively set for Friday the 16th.
January 8, 2009 The chiller is not cooling so I’ve taken the FESEM off-line.
November 21, 2008 The upper SE dector is not working properly, poor image quality. Problem was a piece of dust on the detector plates. I dusted it off and the SE images are fine now.
November 18, 2008 We installed a special fixture on the FESEM stage. You will be able to see the fixture on the chamber scope but it will not interfere with your sample or ability to position the stage. I tested the sccope an it all works fine. Let me know if you have trouble.
October 10. 2008 The FESEM is back up and running. Hitachi repaired the vacuum system and rebuilt the hard drive. The software now runs faster. As always, please save your data to the D drive. Also, I notice today that the EDS is not working properly. I’ll repair and give up an update later.
October 8, 2008 The vacuum system will not stay on. Hitachi will be here tonight to make repair to this problem and to fix the slow computer trouble too.
September 23, 2008 The Hitachi engineer is in town to a)move the SEM 3 feet further from the wall to minimize electric fields and b) install a new hard drive to speed up the computer. Should be done mid afternoon tomorrow.
August 12, 2008 The FESEM is working. I had to reload the Hitachi software and I checked the system operation over but you need to let me know if you observe anything odd. OWen
August 12, 2008 I am still having troubles with the FESEM hard disk. I’ll let you know when it is back working
August 8, 2008 The computer drive is full and has been causing slow operation. I am deleting files to make more room. I’ve setup some program removals and I’m going home. ‘I’ll finish up in the morning. Scope will be back online by 10AM.
July 17, 2008 The stage errors are repaired.
June 26, 2008 The direction of stage movement is reversed. I have made a temporary change that makes the trackball work correctly but the mouse stage movement is still reversed. I am working with Hitachi to resolve the problem. The FESEM can be used while this is being repaired.
March 28, 2008 The Hitachi engineers will perform routine maintenance on both the FESEM and FIB during the week of April 14th. The scopes will be unavailable after noon on the 14th and will be back on line by noon April 18th. Plan your work accordingly.
February 3, 2008 4700 FESEM, reports of inability to stigmate properly. I tested with calibration specimen and it looks fine.
Routine maintenance – The FIB, FESEM and 6400 are all due for factory maintenance. In each case the scope will be down most of a week. I’ll alert you to this schedule.
June 3, 2008 There is an error message that appears when saving images that is the result of the large number of images stored on the hard drive. I’m going to clear them out (tomorrow?) so let me know if you need to save them on another disk.